# This work was supported in part by Hong Kong Innovation and Technology Commission Funding Administrative System II (ITF Ref. No. GHP/110/20GD) and the National Natural Science Foundation of China (62192714, U21B2006).(Corresponding author: Hongwei Liu; Chang Gao.)

### v2

@@ 4 COMPUTATIONAL COMPLEXITY ANALYSIS
The computational complexity of the proposed GLP-MFD algorithm primarily consists of two components: the construction of the observation association graph and the Multi-feature Link Prediction Network, which includes the NFEN layer, the STEF-GAN layer, and the OAJN layer. The complexity of constructing the observation association graph is the function of the observations per frame and the number of frames to be processed. Due to the challenges in detecting low SNR targets, it is often difficult to guarantee detection in every frame. Therefore, each observation needs to attempt associations with those in the subsequent $Q$ frames. When forming graphs from $L$ consecutive frames, for the observations in the $l$-th frame, the maximum velocity constraint is applied for association testing with observations in the set $\{l+1,\ldots,\min\{l+Q,L\}\}$. Given that targets are typically sparse in radar observation space, we explore the computational complexity under the case where only false alarms are present in each frame. For a radar system with $N_{r}$ range cells, $N_{\theta}$ beam positions and $N_{D}$ Doppler channels, when the primary detection threshold $\gamma_{1}$ is set corresponding to the initial false alarm probability $P_{fa,1}$, the average number of required operations $N_{GC}$ for graph construction is as follows[22]:
@@
Let $N_{fa}$ represents the expectation of the false alarm number per frame, then we get $N_{fa}=N_{r}N_{\theta}N_{D}P_{fa,1}$. Consequently, the average computational complexity for graph construction can be simplified to $\mathrm{O}\left(LQN_{fa}^{2}\right)$.
The operations involved in the NFEN layer primarily consist of matrix multiplication and element-wise operations, with its computational complexity being a function of the number of parameters. NFEN is divided into two main categories: the convolutional network and the fully connected networks. The number of convolutional layers is denoted as $N_{\boldsymbol{a}}$. $I_{0}$ and $I_{N_{\boldsymbol{a}}}$ denote the numbers of input and output channels respectively while $I_{p}(p=1,2,\ldots,N_{\boldsymbol{a}}-1)$ denotes the channels from intermediate convolutional operations. $M_{p}(p=1,2,\ldots,N_{\boldsymbol{a}})$ is the size of the convolutional kernel for each layer and the bias term is set to 1. The total number of parameters for the convolutional network is as follows:
@@
$N_{t,d}=2N_{h}+\left(N_{d}-2\right)\left(N_{h}+1\right)N_{h}+\left(N_{h}+1\right)N_{h}$ (20)
$N_{t,s}=2N_{h}+\left(N_{s}-2\right)\left(N_{h}+1\right)N_{h}+\left(N_{h}+1\right)N_{h}$
@@
where each term in the formula corresponds to the parameters of the input layer, hidden layers, and the output layer, respectively. The coefficient of $N_{h}$ in the first term denote the input feature dimension plus a bias which is set to 1. Therefore, the total parameters in NFEN is:
$N_{t,NEFN}=N_{t,d}+N_{t,s}+N_{t,\boldsymbol{i}}+N_{t,\boldsymbol{a}}\text{.}$ (21)
For the STEF-GAN layer, the parameter number of $\mathbf{L}_{E}^{(k)}$ in all $N_{G}$ layers is $2N_{L_{E}}$, where 2 is the feature dimension of $\boldsymbol{e}_{j,i}$ and $N_{L_{E}}$ is the dimension of the enhanced edge feature. The parameter number of $\mathbf{W}_{G}^{(k)}$ is $N_{x_{k}}N_{x_{k+1}}$, where $N_{x_{k}}$ represents the feature dimension of nodes in the $k$-th layer. $\beta_{G}$ is a fully connected layer with $\left(N_{L_{E}}+N_{x_{k+1}}\right)N_{x_{k+1}}$ parameters. The parameter number of $\mathbf{W}_{E}^{(k)}$ in all $N_{G}$ layers is $2N_{W_{E}}$, where $N_{W_{E}}$ is the dimension of the output edge feature in the attention mechanism while $\boldsymbol{h}_{G}^{(k)}$ has $\left(N_{W_{E}}+2N_{x_{k+1}}\right)$ parameters. Therefore, the total number of parameters in STEF-GAN is as follows:
@@
As to the OAJN layers, let $N_{m}$ and $N_{J}$ denote the number of hidden neurons and the number of layers with the bias term set to 1 , respectively. The total number of parameters in OAJN according to (15) is as follows:
$N_{t,OAJN}$ $=\left(2N_{x_{N_{G}}}+2+1+1\right)N_{m}$ (23)
@@
where $M_{\max}$ and $I_{\max}$ represent the maximum kernel size and the maximum number of channels in NEFN, respectively. And $N_{x,\max}$ represents the maximum feature dimension of nodes in STEF-GAN.
@@
In practical engineering applications, the sensor observation space can typically be divided into $N_{k}$ subareas. Tracks are searched and detected independently within each subarea to reduce the computational burden on the system. Thereby the computational complexity of the graph construction decease to $\mathrm{O}\left(LQN_{fa}^{2}/N_{k}\right)$. Additionally, parallelizing the initialization of tracks should enhance computational performance. The number of parameters in the network directly determines its computational complexity, allowing for optimization such as pruning the parameters that have minimal impact on the network. Moreover, GPU parallel processing can be utilized for operations like matrix multiplication to further improve computational efficiency.
5 SIMULATION RESULTS
@@
To show the performance improvement in multi-frame detection achieved through the proposed GLP-MFD, we conduct simulation experiments in a 2-D scenario, but the proposed algorithm can be easily extended to 3-D scenarios. The radar system we simulate is based on the L-band carrier frequency with $2^{\circ}$ azimuth resolution (3dB antenna), 100m range resolution and 32 pulses transmitted at each beam position. As previously mentioned, the primary threshold $\gamma_{1}$ determines the optimal detection performance that the proposed method can achieve. Therefore, to ensure effective detection performance for weak targets while avoiding excessive false alarms that will impose a computational burden on the system, we set $\gamma_{1}$ corresponding to an initial false alarm probability $P_{fa,1}=10^{-3}$. Additionally, it is assumed that false alarms originate from receiver internal noise, which is generally valid for long-range target detection.
@@
When processing multi-frame data using neural networks, the relationship among the second threshold $\gamma_{2}$, the final false alarm probability and the detection probability is complex. Consequently, the false alarm probability and detection probability of the proposed algorithm are calculated using Monte Carlo simulations in this work. Since GLP-MFD transforms plot-level detection into track-level detection, the final false alarm probability $P_{fa,2}$ based on $\gamma_{2}$ is defined as the total number of false tracks divided by the total number of detection units. The detection probability is defined as the total number of correctly detected target tracks divided by the total number of targets appearance in the Monte Carlo experiments.
@@ 5.1 The definition of correct target detection
The proposed GLP-MFD enables direct output of target tracks through the link prediction task in observation association graphs. To evaluate the performance of GLP-MFD, it is necessary to define whether the output track correspond to a correct target track. Due to measurement noise and the dense false alarm environment, it is possible for a false alarm to exhibit a quality superior to that of a true target measurement, which may lead to the inclusion of false alarms in the initiated target track. However, if the existence of false alarms does not affect subsequent target tracking, the initiated target track can still be considered as a correct target track.
In light of the points mentioned above, we employ the Optimal Subpattern Assignment (OSPA) distance to determine whether the output target track has been correctly detected[44]. The OSPA distance is a metric used in target tracking tasks to measure the difference between the set of estimated target states and true target states at a given time by defining a distance measure in the state space. Given that the OSPA distance is essentially a metric for measuring discrepancy between sets, in the context of this work, we use it to quantify the difference between the smoothed confirmed target tracks output from MFLPN and the tracks with correct target states, thereby determining whether the output target tracks are indeed correct. Techniques such as linear regression, quadratic regression, and Kalman filtering can be utilized as smoothing algorithms. The formula for the OSPA distance in the context of our work is as follows:
@@
$\hskip 49.22314pt\left.+\frac{\eta^{\xi}}{N_{\hat{\boldsymbol{X}}}}\left(N_{\hat{\boldsymbol{X}}}-N_{\boldsymbol{X}}\right)\right)^{1/\xi},N_{\boldsymbol{X}}\leq N_{\hat{\boldsymbol{X}}}$
$d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})=d_{\xi}^{(\eta)}(\hat{\boldsymbol{X}},\boldsymbol{X})\hskip 42.10995pt,N_{\boldsymbol{X}}>N_{\hat{\boldsymbol{X}}}$
@@
$\xi$ affects the deviation between the estimation and the true value. As $\xi$ increases, the weight of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ relative to the position error enhances. $\eta$ influences the discrepancy in the number of plots between smoothed tracks and correct tracks. As $\eta$ increases, the weight of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ with respect to the mismatch in the number of plots enhances. In the context of our work, we focus on the impact of false alarms on the quality of target tracks, with greater emphasis placed on position errors. Considering practicality and the smoothness of distance curves, we set $\xi=2$. Additionally, the value of $\eta$ should be relatively small. After evaluating the position measurement errors in terms of the variance of measurement noise, the value of $\eta$ determined based on the position measurement errors can be regarded as a small value[45]. Given that the range of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ is $[0,\eta]$, when the OSPA distance between smoothed tracks and tracks with correct target states meets the following condition:
@@
(a) Detection result of the primary threshold
(b) Processing result of the proposed algorithm
Figure 6: Results for one exemplar run

### v3

@@ 4 COMPUTATIONAL COMPLEXITY ANALYSIS
The computational complexity of the proposed GLP-MFD algorithm primarily consists of two components: the construction of the observation association graph and the processing of MFLPN, whose network module includes NFEN, STEF-GAT, and OAJN. The complexity of constructing the observation association graph is the function of the observations per frame and the number of frames to be processed. Due to the challenges in detecting low SNR targets, it is often difficult to guarantee detection in every frame. Therefore, each observation needs to attempt associations with those in the subsequent $Q$ frames. When forming graphs from $L$ consecutive frames, for the observations in the $l$-th frame, the maximum velocity constraint is applied for association testing with observations in the set $\{l+1,\ldots,\min\{l+Q,L\}\}$. Given that targets are typically sparse in the observation space, we explore the computational complexity under the case where only false alarms are present in each frame. For a sensor system with $N_{r}$ range cells, $N_{\theta}$ beam positions and $N_{D}$ Doppler channels, when the primary detection threshold $\gamma_{1}$ is set corresponding to the initial false alarm probability $P_{fa,1}$, the average number $N_{GC}$ of required operations for graph construction is as follows [23]:
@@
Let $N_{fa}$ represent the expectation of the false alarm number per frame, then we get $N_{fa}=N_{r}N_{\theta}N_{D}P_{fa,1}$. Consequently, the average computational complexity for graph construction can be simplified to $\mathrm{O}\left(LQN_{fa}^{2}\right)$.
The operations involved in NFEN primarily consist of matrix multiplication and element-wise operations. Its computational complexity is a function of the number of parameters. NFEN is divided into two main categories: the convolutional network and the fully connected networks. The number of convolutional layers is denoted as $N_{\boldsymbol{a}}$. $I_{0}$ and $I_{N_{\boldsymbol{a}}}$ denote the numbers of input and output channels respectively while $I_{p}(p=1,2,\ldots,N_{\boldsymbol{a}}-1)$ denotes the channels from intermediate convolutional operations. $M_{p}(p=1,2,\ldots,N_{\boldsymbol{a}})$ is the size of the convolutional kernel for each layer and the bias term is set to 1. The total number of parameters for the convolutional network is as follows:
@@
$N_{t,d}$ $=2N_{h}+\left(N_{d}-2\right)\left(N_{h}+1\right)N_{h}+\left(N_{h}+1\right)N_{h}\text{,}$ (21)
$N_{t,s}$ $=2N_{h}+\left(N_{s}-2\right)\left(N_{h}+1\right)N_{h}+\left(N_{h}+1\right)N_{h}\text{,}$
@@
where each term in the formula corresponds to the parameters of the input layer, hidden layers, and the output layer, respectively. The coefficient of $N_{h}$ in the first term denotes the input feature dimension plus a bias which is set to 1. Therefore, the total parameters in NFEN is:
$N_{t,NFEN}=N_{t,d}+N_{t,s}+N_{t,\boldsymbol{i}}+N_{t,\boldsymbol{a}}\text{.}$ (22)
For STEF-GAT, the parameter number of $\mathbf{L}_{E}^{(k)}$ in all $N_{G}$ layers is $2N_{L_{E}}$, where 2 is the feature dimension of $\boldsymbol{e}_{j,i}$ and $N_{L_{E}}$ is the dimension of the enhanced edge feature. The parameter number of $\mathbf{W}_{G}^{(k)}$ is $N_{x_{k}}N_{x_{k+1}}$, where $N_{x_{k}}$ represents the feature dimension of nodes in the $k$-th layer. $\beta_{G}$ is a fully connected layer with $\left(N_{L_{E}}+N_{x_{k+1}}\right)N_{x_{k+1}}$ parameters. The parameter number of $\mathbf{W}_{E}^{(k)}$ in all $N_{G}$ layers is $2N_{W_{E}}$, where $N_{W_{E}}$ is the dimension of the output edge feature in the attention mechanism. The parameter number of $\boldsymbol{a}_{G}^{(k)}$ is $\left(N_{W_{E}}+2N_{x_{k+1}}\right)$. Therefore, the total number of parameters in STEF-GAT is as follows:
@@
As to OAJN, let $N_{m}$ and $N_{J}$ denote the number of hidden neurons and the number of layers with the bias term set to 1 , respectively. The total number of parameters in OAJN according to (16) is as follows:
$N_{t,OAJN}$ $=\left(2+1+1\right)N_{m}+\left(2N_{x_{N_{G}}}+N_{m}+1\right)N_{m}$ (24)
@@
where $M_{\max}$ and $I_{\max}$ represent the maximum kernel size and the maximum number of channels in NFEN, respectively. And $N_{x,\max}$ represents the maximum feature dimension of nodes in STEF-GAT.
@@
In practical engineering applications, the entire observation space can typically be divided into $N_{\tau}$ subareas. Tracks are searched and detected within each subarea to reduce the computational burden on the system. Therefore, the computational complexity of the graph construction will decrease to $\mathrm{O}\left(LQN_{fa}^{2}/N_{\tau}\right)$. In addition, parallelizing the initiation of tracks in subareas should enhance computational performance. The number of parameters in the network determines its computational complexity, allowing for optimization such as pruning the parameters that have minimal impact on the network. Moreover, GPU parallel processing can be utilized for operations like matrix multiplication to further improve computational efficiency of the proposed algorithm.
5 NUMERICAL RESULTS
@@
In this section, numerical experiments are conducted to show the performance improvement in multi-frame detection achieved by the proposed GLP-MFD. The primary threshold $\gamma_{1}$ determines the optimal detection performance that the proposed algorithm can achieve. Therefore, to ensure effective detection performance for weak targets while avoiding excessive false alarms that will impose a computational burden on the system, we set $\gamma_{1}$ corresponding to an initial false alarm probability $P_{fa,1}=10^{-3}$. Additionally, it is assumed that false alarms originate from internal receiver noise, which is generally valid for long-range target detection.
@@
When processing multi-frame data using neural networks, the relationship among $\gamma_{2}$, the final false alarm probability, and the detection probability is complex. Consequently, the final false alarm probability and detection probability of the proposed algorithm are calculated using Monte Carlo experiments in this work. Since GLP-MFD transforms plot-level detection into track-level detection, the final false alarm probability $P_{fa,2}$ based on $\gamma_{2}$ is defined as the total number of false tracks divided by the total number of detection cells. The detection probability is defined as the total number of correctly detected target tracks divided by the total number of target appearances in the Monte Carlo experiments.
Figure 7: Confusion matrices for different SNRs.
5.1 Network performance analysis
The proposed MFLPN performs a link prediction task in observation association graphs and is implemented using the PyTorch framework [52]. The dataset is simulated based on an L-band radar with $2^{\circ}$ azimuth resolution (3dB antenna), 100 m range resolution and 32 pulses transmitted at each beam position. The graph used for training is constructed by multi-frame observations that exceed the primary threshold, based on the method described in section 3.3.1. The hardware configuration is NVIDIA GeForce RTX 4090 GPU. We set the batch size to 32 and use the Adam optimizer for training the network [53]. The total number of training epochs is 800 and the initial learning rate is set to 0.01, which is reduced to 1/10 of its previous value every 200 epochs. To show the key hyperparameters in the proposed network, we provide the detailed network structure as shown in Fig 6. For better readability, we take the link prediction between node $i$ and node $j$ in a graph as an example. The description “Conv(4,1,3$\times$3)” signifies a convolutional layer with 4 input channels, 1 output channel and a 3$\times$3 kernel size. “p=(5,0)” refers to a padding of 5 in height and 0 in width, with a circular padding mode. If description “p” is not provided, no padding operation is performed. “s=(2,1)” refers to a stride of 2 in height and 1 in width. If “s” is not provided, the default stride is 1 in both height and width. “Linear(1,4)” denotes a fully connected layer with an input dimension of 1 and an output dimension of 4. “MP(16,32)” denotes a message passing layer with an input dimension of 16 and an output dimension of 32. The meaning of the remaining descriptions can be inferred by analogy.
Link prediction can be regarded as a classification task for edges in the graph. In this work, the edges are classified into three categories: false alarm to false alarm (F-F), target to false alarm (T-F), and target to target (T-T). To demonstrate the network performance, we present the confusion matrices under different SNR conditions, as shown in Fig. 7. It can be observed that targets with higher SNR, due to their more distinct target characteristic, result in better classification performance. In addition, by observing the network accuracy for each structure, we conduct the ablation study to validate the effectiveness of each individual module in the proposed network. Moreover, to further demonstrate the advantage of message passing in STEF-GAT, we choose the Graph Convolutional Network (GCN) for comparison [54]. GCN is a classical method in graph neural networks and we replace STEF-GAT with GCN in the full network. The network accuracy for various SNRs in the ablation study is shown in Table 1. The result indicates that the designed NFEN, STEF-GAT and OAJN all contribute positively to the overall network performance. NFEN has the ability to effectively pre-extract node features, aiding the network training process. OAJN facilitates better utilization of the extracted features to accomplish the required task. STEF-GAT, as the most crucial module in the network, is capable of learning the spatio-temporal coupling between multiple frames, further extracting node features with multi-frame fusion. It maximally supports the proposed algorithm to perform integrated track search and track detection in observation association graphs. Since the message passing in GCN fails to incorporate the attention mechanism and does not design edge features based on spatio-temporal coupling information, the contribution of GCN is far inferior to that of STEF-GAT. This comparison further highlights the superiority of the designed message passing in STEF-GAT.
Table 1: Ablation study result.
Structure 7dB 8dB 9dB 10dB 11dB 12dB
   NFEN 68.19% 70.30% 74.24% 77.23% 80.70% 82.37%
   STEF-GAT 71.48% 76.50% 80.52% 85.60% 89.04% 90.68%
   OAJN 64.71% 67.32% 70.65% 73.94% 77.41% 79.36%
   NFEN + STEF-GAT 81.18% 85.02% 88.63% 92.26% 94.02% 95.32%
   NFEN + OAJN 72.94% 75.24% 78.14% 81.04% 84.13% 85.73%
   STEF-GAT + OAJN 77.71% 82.47% 86.19% 89.60% 92.46% 93.77%
   NFEN + GCN + OAJN 74.36% 78.20% 80.46% 84.26% 85.97% 87.22%
   NFEN + STEF-GAT + OAJN 87.60% 90.84% 93.33% 95.51% 96.73% 97.44%
Table 2: Network accuracy with different layers in STEF-GAT.
Layer number 1 2 3 4 5
Accuracy 89.55% 91.15% 93.58% 91.64% 91.87%
Table 3: Runtimes of the network.
Node Number $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$
NFEN + STEF-GAT + OAJN GPU 0.0069 s 0.0082 s 0.0172 s 0.1644 s
CPU 0.0085 s 0.0443 s 0.2439 s 1.5588 s
NFEN + GCN + OAJN GPU 0.0053 s 0.0058 s 0.0108 s 0.1322 s
CPU 0.0054 s 0.0292 s 0.1626 s 1.0801 s
The configuration of hyperparameters should take into account both practical requirements and physical significance. Given that the radial velocity of the moving target varies across multiple frames, it is crucial to ensure that the receptive field of the convolutional neural network in NFEN can cover the changes in the target Doppler channels. The fully connected networks in NFEN aim to map input features to a higher dimensional space and the purpose of OAJN is to better perform link prediction based on the features extracted by previous modules. The multi-layer message passing in STEF-GAT implies the fusion of multi-frame information. However, it is difficult to ensure successful detection of weak targets in each frame. Therefore, it is necessary to balance the number of frames fused for message passing of multi-frame target observations with the increased training difficulty and computational cost brought by excessively deep layers. Methods such as grid search, random search, and Bayesian optimization can be used to jointly tune the hyperparameters [55]. When the hyperparameter search space is large, Bayesian optimization can significantly reduce the search iterations and more efficiently identify a hyperparameter configuration that is close to the optimum. Bayesian optimization commonly involves the following steps. First, an objective to be optimized and the input hyperparameter space are defined. Next, a surrogate model like Gaussian process is selected to fit the objective function. Then, an acquisition function like expected improvement is used to sample the hyperparameter space and evaluate the objective. Finally, the surrogate model is iteratively updated based on the evaluation results. When tuning the hyperparameters of the proposed network, the network accuracy can be chosen as the objective to be optimized, and the hyperparameter search space can be defined based on each module. For example, the search range for the number of convolutional layers in NFEN can be set between 1 and 8, while the range for the number of layers in STEF-GAT and OAJN can be set between 1 and 5. The number of hidden neurons can range from 8 to 64, while the learning rate and batch size can range from $10^{-1}$ to $10^{-5}$ and 16 to 64, respectively. The tool scikit-optimize can be used for hyperparameter optimization of the proposed network. In addition, to further demonstrate the impact of the number of layers in the key model STEF-GAT on network performance, we provide the average network accuracy across various SNRs for different layer configurations during message passing, as shown in Table 2. The best performance occurs with 3 layers of message passing. When the number of layers is small, the accuracy may decline due to insufficient fusion of multi-frame information. Conversely, when the number of layers is large, the accuracy may degrade due to increased training difficulty, such as overfitting.
To validate the time cost of the proposed network, we calculate the average runtime for an observation association graph using the Monte Carlo experiment. The runtime of a graph is directly related to its scale. As the graph scale increases, the time cost will correspondingly rise. The graph scale is closely related to factors such as the size of the divided subareas, the initial false alarm probability for the primary detection threshold, the number of processing frames, and the number of subareas processed in parallel. For a more intuitive demonstration, we reflect the graph scale by the number of nodes in the graph. For example, in a subarea with $10^{5}$ detection cells, where the initial false alarm probability is $10^{-3}$, with $5$ processing frames and $20$ subareas processed in parallel, the number of nodes in the graph is approximately $10^{4}$. As shown in Table 3, we provide the runtime for different graph scales on Intel(R) Xeon(R) Platinum 8370C CPU and NVIDIA GeForce RTX 4090 GPU. To further validate the computational efficiency of the key module STEF-GAT, we replace it with GCN in the full network. Therefore, we test the runtime of the network consisting of NFEN, STEF-GAT, and OAJN, as well as the network consisting of NFEN, GCN, and OAJN. It can be seen that for a network including a large number of matrix multiplications and element-wise operations, GPU has a significant advantage in computational efficiency compared to CPU. When the graph scale is relatively large, this advantage becomes more prominent. Since GCN has fewer parameters, it takes less time to run. The runtime of the network including STEF-GAT is on the same order of magnitude as the network including GCN, indicating that STEF-GAT does not bring a notable computational cost. However, as shown in Table 1, the contribution of STEF-GAT to the improvement in network accuracy is significant compared to GCN.
@@ 5.2 The definition of correct target detection
The proposed GLP-MFD enables direct output of target tracks through the link prediction task in observation association graphs. To evaluate the performance of GLP-MFD, it is necessary to define whether the output track correspond to a correct target track. Due to measurement noise and the dense false alarm environment, it is possible for a false alarm to exhibit a quality superior to that of a true target measurement, which may lead to the inclusion of false alarms in the initiated target track. However, if the existence of false alarms does not affect subsequent target tracking, the initiated target track can still be considered as a correct target track. In light of the points mentioned above, we employ the optimal subpattern assignment (OSPA) distance [56] to determine whether the output target track has been correctly detected. The OSPA distance is a metric used in target tracking tasks to measure the difference between the set of estimated target states and true target states at a given time, by defining a distance measure in the state space. Given that the OSPA distance is essentially a metric for measuring the discrepancy between sets, in the context of this work, we use it to quantify the difference between the smoothed confirmed target tracks output from MFLPN and the tracks with correct target states. In this manner, we determine whether the output target tracks are indeed correct. Techniques such as linear regression, quadratic regression, and Kalman filtering can be utilized as smoothing algorithms. The formula for the OSPA distance in the context of our work is as follows:
@@
$\hskip 49.22314pt\left.+\frac{\eta^{\xi}}{N_{\hat{\boldsymbol{X}}}}\left(N_{\hat{\boldsymbol{X}}}-N_{\boldsymbol{X}}\right)\right)^{1/\xi},N_{\boldsymbol{X}}\leq N_{\hat{\boldsymbol{X}}}\text{,}$
$d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})=d_{\xi}^{(\eta)}(\hat{\boldsymbol{X}},\boldsymbol{X})\hskip 42.10995pt,N_{\boldsymbol{X}}>N_{\hat{\boldsymbol{X}}}\text{,}$
@@
$\xi$ affects the deviation between the estimation and the true value. As $\xi$ increases, the weight of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ relative to the position error enhances. $\eta$ influences the discrepancy in the number of plots between smoothed tracks and correct tracks. As $\eta$ increases, the weight of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ with respect to the mismatch in the number of plots enhances. In the context of our work, we focus on the impact of false alarms on the quality of target tracks, with greater emphasis placed on position errors. Considering practicality and the smoothness of distance curves, we set $\xi=2$ [57]. Moreover, the value of $\eta$ should be relatively small. After evaluating the position measurement errors in terms of the variance of measurement noise, the value of $\eta$ determined based on the position measurement errors can be regarded as a small value [57]. Given that the range of $d_{\xi}^{(\eta)}(\boldsymbol{X},\hat{\boldsymbol{X}})$ is $[0,\eta]$, when the OSPA distance between smoothed tracks and tracks with correct target states meets the following condition:
@@
(a) Detection result of the primary threshold.
(b) Processing result of the proposed algorithm.
Figure 9: Results for one exemplar run.